p { margin-bottom: 0.25cm; line-height: 120%; }a:link { }

The position of a particle moving along X-axis is given by x=At+4Bt3.What is its velocity at t=5s?

Grade:11

1 Answers

Onkar Joshi
6 years ago
Position of vector=At+Bt^3Velocity is given by=dx/dtVelocity=d(At+Bt^3)/dt=A+12Bt^2So velocity at 5s= A+12*(5)^2B=A+300B